OpenStudy (anonymous):

Please help number nine says the position of an object at time t moving along a line is given by the function s(t)=-16t^2+128t find the average velocity of the object over the following intervals [1,4] [1,3] [1,2] [1,1+h]

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

If can just help with first one I could use that as sample I am looking at book example but not understanding it

OpenStudy (anonymous):

The average velocity over an interval is given by the slope of the secant line over the interval: \[v_{avg}=\frac{s(b)-s(a)}{b-a}\] where the interval is \([a,b]\).

OpenStudy (anonymous):

With \(s(t)=-16t^2+128t\) and the first interval \([1,4]\), you have \[v_{avg}=\frac{\left(-16(4)^2+128(4)\right)-\left(-16(1)^2+128(1)\right)}{4-1}\]

OpenStudy (anonymous):

Oooo ok

OpenStudy (anonymous):

That makes more sense now thanks :) so for this one my answer would be 48 for first one :)

OpenStudy (anonymous):

Right. You're welcome!